Numbers 6,7, and 8 please
A) (24,-2) 6)[y=5x + 7 y=8x + 6 A) infinitely many solutions C) no solution (1 26) 3 3 26 1 Solve the system of equations. 7) y=4x+1 3y-9x = 15 A) (17, 4) C) l(x, y)ly B) (4, 17) D) ø 4x +1) 8) 3x +8y -2 2x+5y =-7 A) (-46, 17) B)(-46-17 C) (17,-46) D) (-17,46)
Answer #1

6)Given equations are v 5x7 У 8x +6 2 From equ 1 and equ2 8x-5x = 7-6 Sub x value in equi +7 5 +21 26 1 26 So the correct option is B 7)Given equations are У 4х +1 ! 31-9x 15 2 From equ 1 and equ2 3(4x+1)-9-15 12x + 3-9x 15 3x-15-3 3.x-12 Sub x value in equi У 4(4)+1 = 16+1 17 (x.y)-(4.17) So the correct option is B8) 3x +8y =-2 1 2x+5y... equ1*2 -equ2*3 6x +16y-4 6x+15y =-21 Subtract these equations 6x +163-6x-15y=-4-(-21) 16y-5y--4+21 V-17 Sub y value in equl 3x+(8)(17)--2 3x 136 : :. 3x--138 (x,y)-(-46,17) So the correct option is A